# Biomedical Entity Linking as Multiple Choice Question Answering

### v1

@@ Abstract
[…]
@@ 1.   Introduction
Existing methods can be mainly categorized into two types. One is the discriminative methods Sung et al. (2020); Lai et al. (2021); Liu et al. (2021), which employed BERT-based models to encode mentions and entities into the same embedding space and disambiguated mentions by nearest neighbors, or further applied a cross-encoder to rerank top candidates by capturing fine-grained mention-entity interactions Angell et al. (2021); Zhu et al. (2021); Xu et al. (2023). Another one is the generative methods Yan et al. (2020); Yuan et al. (2022a, b) that directly generated linked entities using text-to-text pre-trained language models, such as BART Lewis et al. (2020), thereby circumventing the need for negative sample mining.
@@
To tackle the aforementioned challenges, we present a novel model called BioELQA, which aims at capturing the fine-grained interactions between entities and enhancing the generality of long-tailed entities. We achieve this by reframing BioEL as a Multiple Choice Question Answering (MCQA) task. Given a mention, we first use a bi-encoder retriever to efficiently retrieve top-$N$ candidate entities from the ontology as the “answer options”. Then the mention and its answer options are all fed to a generator as a multiple choice prompt, with each answer associated with a symbol (e.g., "A", "B", "C"). The prompt is structured so that the generator outputs the symbol associated with its chosen answer option. This framework enables explicit comparison and contrast among different answer options, effectively modeling the mention-entity interactions and entity-entity interactions. Besides, by directly generating answer symbols instead of entity names, we can avoid generating entities that do not exist in the ontology and reduce reliance on normalization strategies Robinson et al. (2022). Inspired by the recent success of retrieval-augmented model Khandelwal et al. (2020), we further leverage a $k$NN module to retrieve semantically similar instances from the training data to formulate a retrieval-enhanced prompt, which empowers the model to reference similar contexts as cues for prediction, to improve the robustness and generalization for long-tailed entities. We experimentally validate the effectiveness and superiority of BioELQA as it achieves state-of-the-art performance on several benchmark datasets. The source code and curated prompts will be made publicly accessible after the review.
@@ 3.1.   Experimental Setup
Baselines. We compare the proposed BioELQA against previous state-of-the-art BioEL methods, which are classified into three categories: 1) discriminative-based methods that utilize bi-encoders or extend their capabilities with cross-encoders to retrieve relevant entities, including BioSyn Sung et al. (2020), ResCNN Lai et al. (2021), SapBERT Liu et al. (2021), Clustering-based Angell et al. (2021) and Prompt-BioEL Xu et al. (2023); 2) generative-based methods that directly generate the linked entities, including GenBioEL Yuan et al. (2022b) and BioBART Yuan et al. (2022a); 3) LLM-based methods that have demonstrated impressive capabilities across various tasks, especially in the biomedical domain, with simple instructions Jahan et al. (2023), including GPT3.5, [^1] https://platform.openai.com/docs/models/gpt-3-5 PaLM-2 Anil et al. (2023), Claude-2, [^2] https://www.anthropic.com/index/claude-2 and LLaMA-2-13b Touvron et al. (2023). Note that the results of LLM-based methods are taken from Jahan et al. (2023).
@@ 3.4.   Case Study
Mention Retrieved Instances
Prediction
seizure medication (seizure meds, seizure management), (migraine treatments, migraine prophylaxis)
($Q$) seizure management
@@
sketchy (sketchy, incomplete), (blurry, blurring of visual image)
($Q$) incomplete
@@
rhodiola (rhodiola, family crassulaceae), (broccoli, brassica oleracea)
($Q$) family crassulaceae
@@
lose weight (losing weight, weight loss)
($Q$) weight loss
@@
Mention Context
Prediction
suppositories for some personal reasons i can’t use the [E1] suppositories [/E1]
($Q$) suppository
@@
salt i guess maybe some [E1] salt [/E1] residue gets on the insertion site
($Q$) salt water
@@
sublingual maxalt comes in a [E1] sublingual [/E1] under the tongue disolve version
($Q$) sublingual
@@
gum can’t really see it but there’s a dark line underneath my front teeth what seems like around my [E1] gum [/E1] shown above
($Q$) gum
@@ 5.   Bibliographical References
  • Lewis et al. (2020) Mike Lewis, Yinhan Liu, Naman Goyal, Marjan Ghazvininejad, Abdelrahman Mohamed, Omer Levy, Veselin Stoyanov, and Luke Zettlemoyer. 2020. BART: Denoising Sequence-to-Sequence Pre-training for Natural Language Generation, Translation, and Comprehension. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 7871–7880.
@@
  • Lin et al. (2024) Zhenxi Lin, Ziheng Zhang, Xian Wu, and Yefeng Zheng. 2024. Improving Biomedical Entity Linking with Retrieval-enhanced Learning. In ICASSP 2024 - 2024 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP).

### v2

@@ Abstract
Zhenxi Lin1, Ziheng Zhang1, Xian Wu, Yefeng Zheng
@@ 1.   Introduction
Existing methods can be mainly categorized into two types. One is the discriminative methods Sung et al. (2020); Lai et al. (2021); Liu et al. (2021); Wu et al. (2023), which employed BERT-based models to encode mentions and entities into the same embedding space and disambiguated mentions by nearest neighbors, or further applied a cross-encoder to rerank top candidates by capturing fine-grained mention-entity interactions Angell et al. (2021); Zhu et al. (2021); Xu et al. (2023). Another one is the generative methods Yan et al. (2020); Yuan et al. (2022a, b) that directly generated linked entities using text-to-text pre-trained language models, such as BART Lewis et al. (2020), thereby circumventing the need for negative sample mining.
@@
To tackle the aforementioned challenges, we present a novel model called BioELQA, which aims at capturing the fine-grained interactions between entities and enhancing the generality of long-tailed entities. We achieve this by reframing BioEL as a Multiple Choice Question Answering (MCQA) task. Given a mention, we first use a bi-encoder retriever to efficiently retrieve top-$N$ candidate entities from the ontology as the “answer options”. Then the mention and its answer options are all fed to a generator as a multiple choice prompt, with each answer associated with a symbol (e.g., "A", "B", "C"). The prompt is structured so that the generator outputs the symbol associated with its chosen answer option. This framework enables explicit comparison and contrast among different answer options, effectively modeling the mention-entity interactions and entity-entity interactions. Besides, by directly generating answer symbols instead of entity names, we can avoid generating entities that do not exist in the ontology and reduce reliance on normalization strategies Robinson et al. (2022). Inspired by the recent success of retrieval-augmented model Khandelwal et al. (2020), we further leverage a $k$NN module to retrieve semantically similar instances from the training data to formulate a retrieval-enhanced prompt, which empowers the model to reference similar contexts as cues for prediction, to improve the robustness and generalization for long-tailed entities. We experimentally validate the effectiveness and superiority of BioELQA as it achieves state-of-the-art performance on several benchmark datasets. The source code and curated prompts are publicly accessible at https://github.com/lzxlin/BioELQA.
@@ 3.1.   Experimental Setup
Baselines. We compare the proposed BioELQA against previous state-of-the-art BioEL methods, which are classified into three categories: 1) discriminative-based methods that utilize bi-encoders or extend their capabilities with cross-encoders to retrieve relevant entities, including BioSyn Sung et al. (2020), ResCNN Lai et al. (2021), SapBERT Liu et al. (2021), Clustering-based Angell et al. (2021) and Prompt-BioEL Xu et al. (2023); 2) generative-based methods that directly generate the linked entities, including GenBioEL Yuan et al. (2022b) and BioBART Yuan et al. (2022a); 3) LLM-based methods that have demonstrated impressive capabilities across various tasks, especially in the biomedical domain, with simple instructions Jahan et al. (2023), including GPT3.5, [^1] https://platform.openai.com/docs/models/gpt-3-5 PaLM-2 Anil et al. (2023), Claude-2, [^2] https://www.anthropic.com/index/claude-2 and LLaMA-2-7b/13b Touvron et al. (2023). Note that the results of LLM-based methods are all taken from Jahan et al. (2023).
@@ 3.4.   Case Study
Mention Retrieved Instances Prediction
seizure medication (seizure meds, seizure management), (migraine treatments, migraine prophylaxis) ($Q$) seizure management
@@
sketchy (sketchy, incomplete), (blurry, blurring of visual image) ($Q$) incomplete
@@
rhodiola (rhodiola, family crassulaceae), (broccoli, brassica oleracea) ($Q$) family crassulaceae
@@
lose weight (losing weight, weight loss) ($Q$) weight loss
@@
Mention Context Prediction
suppositories for some personal reasons i can’t use the [E1] suppositories [/E1] ($Q$) suppository
@@
salt i guess maybe some [E1] salt [/E1] residue gets on the insertion site ($Q$) salt water
@@
sublingual maxalt comes in a [E1] sublingual [/E1] under the tongue disolve version ($Q$) sublingual
@@
gum can’t really see it but there’s a dark line underneath my front teeth what seems like around my [E1] gum [/E1] shown above ($Q$) gum
@@ 5.   Bibliographical References
  • Lewis et al. (2020) Mike Lewis, Yinhan Liu, Naman Goyal, Marjan Ghazvininejad, Abdelrahman Mohamed, Omer Levy, Veselin Stoyanov, and Luke Zettlemoyer. 2020. BART: Denoising Sequence-to-Sequence Pre-training for Natural Language Generation, Translation, and Comprehension. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics.
@@
  • Lin et al. (2024) Zhenxi Lin, Ziheng Zhang, Xian Wu, and Yefeng Zheng. 2024. Improving biomedical entity linking with retrieval-enhanced learning. In ICASSP 2024 - 2024 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP), pages 11461–11465.
@@
  • Wu et al. (2023) Taiqiang Wu, Xingyu Bai, Weigang Guo, Weijie Liu, Siheng Li, and Yujiu Yang. 2023. Modeling Fine-grained Information via Knowledge-aware Hierarchical Graph for Zero-shot Entity Retrieval. In Proceedings of the Sixteenth ACM International Conference on Web Search and Data Mining, pages 1021–1029.